# GeoScaler: Geometry and Rendering-Aware Downsampling of 3D Mesh Textures

### 28 Nov 2023 v1

@@ 5 Conclusions
We presented a method of downsampling texture maps that is superior to existing methods using a unified representation of texture maps and the geometries of 3D meshes that they are painted on. This framework is also applicable for downsampling other feature maps associated with meshes, such as roughness and specularity.
@@
GeoScaler-based downsampling of textures is intended to enable the deployment of compact graphical assets in applications for devices having low computation and memory budgets such as smartphones and VR headsets while maintaining high visual quality. It also facilitates remastering high-quality graphical assets for video games on devices with lower rendering capabilities.
Supplementary Material
6 More Results
Visual results of meshes (avocado, banjoman, and shoe2) included in the full paper are shown in Fig. 7 for the sake of completeness. All mesh .obj files with the original, bicubic downsampled, and GeoScaler downsampled texture maps are also available in the submitted supplementary material.
Figure 7: Qualitative results comparing GeoScaler with bicubic downsampling for 4x and 8x downsampling. For each mesh result, the top row is the Ground Truth texture, the middle row shows results from GeoScaler, and the bottom row shows results using Bicubic.
A more detailed mesh-wise ablation is shown in Fig. 8 and Fig. 9 for 4x and 8x downsampling respectively. Fig. 8 and Fig. 9 also show the quality of images rendered using textures generated at lower resolutions by Metashape directly while performing the 3D reconstruction process. These results suggest that downsampling textures after reconstructing 3D scenes at higher resolutions (using tools such as MetaShape) can provide higher-quality renders than generating low-resolution textures directly. In these tables, it is also evident that Lanczos resampling and bicubic resampling are extremely similar quantitatively. Given this observation and the widespread popularity of bicubic resampling, the results of Lanczos resampling were omitted from the full paper for the sake of brevity. Qualitative results of the ablations for stew and bevo are also shown in Fig. 11.
Figure 8: Mesh-wise Ablations for 4x Downsampling. GeoScaler outperforms traditional methods and also performs better than generating textures natively at lower resolutions
Figure 9: Mesh-wise Ablations for 8x Downsampling. GeoScaler outperforms traditional methods and also performs better than generating textures natively at lower resolutions
Figure 10: Comparing Texture maps generated by bicubic downsampling and GeoScaler. Textures generated by GeoScaler have larger areas assigned to regions with high detail density (as shown by red boxes) and are more sharpened to compensate for the downsampling
Figure 11: Visual Results of Ablation Study. The independent contributions of the GeoCoding Module and the UVWarping module are illustrated. GeoCoding stitches artifacts arising due to seams in texture maps while UVWarping improves the details in the rendered images
7 A Closer Look at the Generated Textures
Instead of analyzing rendered images, we study the texture maps themselves. The difference between texture maps downsampled by bicubic and GeoScaler is depicted in Fig. 10, which contains crops of 8x downsampled texture maps of avocado, clothdolls and stew meshes. It can be observed that the texture maps generated by GeoScaler look over-sharpened. However, it is important to note that texture maps are first wrapped onto surfaces of meshes (whose surface areas remain unchanged throughout the downsampling process) before views are rendered using them. As surface areas of meshes remain unchanged, the downsampled textures have to be stretched/upsampled to cover them. The over-sharpening and the upsampling compensate for each other’s detrimental distortion and produce a rendered result that resembles that of the original texture. GeoScaler is capable of estimating the degrees of (over-)sharpening needed to apply on different regions of texture maps based on ratios of areas of faces on mesh surfaces and their corresponding UV planes.
In Fig. 10 the effect of the UVWarper module is also visible. The regions marked by the red boxes occupy a larger area in texture maps produced by GeoScaler compared to the area they cover in the texture maps produced using bicubic downsampling. Without any additional semantic guidance/supervision, the UVWarper is capable of recognizing areas of important textural information and consequently assigning larger areas to them in the generated texture map. The warping of texture maps reduces the loss of information occurring during the downsampling of the textures and enhances the perceptual quality of rendered images.

### 20 Mar 2024 v2

@@ 5 Conclusions
We presented a method of downsampling texture maps that is superior to existing methods using a unified representation of texture maps and the geometries of 3D meshes that they are painted on. This framework is also applicable for downsampling other feature maps associated with meshes, such as roughness and specularity. GeoScaler-based downsampling of textures is intended to enable the deployment of compact graphical assets in applications for devices having low computation and memory budgets such as smartphones and VR headsets while maintaining high visual quality. It also facilitates remastering high-quality graphical assets for video games on devices with lower rendering capabilities.
While direct potential negative impacts of this work are not anticipated, the method has a few limitations. Firstly, this method cannot be directly used to generate texture maps from images captured in the wild as it relies on the presence of a high-resolution texture map. Consequently, this method is also unable to rectify and would replicate any errors or artifacts present in the original texture. Secondly, while a per-mesh optimization framework enables content-adaptive downsampling and significantly enhanced quality, it requires much more processing time than the existing methods that are practically instantaneous. These limitations pave the way for interesting future work on this topic.